Let Z be the standard normal random variable. Find P(Z > 2.15)

P(0.25< Z < 1.77)


1
Expert's answer
2021-11-29T13:57:38-0500

"P(Z>2.15) = 1 -P(Z<2.15) \\
=1 -0.9842 \\
= 0.0158 \\
P(0.25<Z<1.77) = P(Z<1.77) -P(Z<0.25) \\
= 0.9616 -0.5987 \\
= 0.3629"